OpenStudy (anonymous):

What is the value of t? -12.8= 4t t =__

OpenStudy (anonymous):

@jordanloveangel

OpenStudy (anonymous):

4?

OpenStudy (zpupster):

divide both sides by 4 that will give you t

OpenStudy (jordanloveangel):

one sec

OpenStudy (anonymous):

no wait

OpenStudy (anonymous):

i didn't finsh so sorry

OpenStudy (anonymous):

-3.2

OpenStudy (jordanloveangel):

no

OpenStudy (jordanloveangel):

ur answer is wrong

OpenStudy (anonymous):

?

OpenStudy (jordanloveangel):

-12.8=4t Simple and best practice solution for -12.8=4t equation. Check how easy it is, and learn it for the future. Our solution is simple, and easy to understand, so don`t hesitate to use it as a solution of your homework. If it's not what You are looking for type in the equation solver your own equation and let us solve it. Equation: Solution for -12.8=4t equation: Simplifying -12.8 = 4t Solving -12.8 = 4t Solving for variable 't'. Move all terms containing t to the left, all other terms to the right. Add '-4t' to each side of the equation. -12.8 + -4t = 4t + -4t Combine like terms: 4t + -4t = 0 -12.8 + -4t = 0 Add '12.8' to each side of the equation. -12.8 + 12.8 + -4t = 0 + 12.8 Combine like terms: -12.8 + 12.8 = 0.0 0.0 + -4t = 0 + 12.8 -4t = 0 + 12.8 Combine like terms: 0 + 12.8 = 12.8 -4t = 12.8 Divide each side by '-4'. t = -3.2 Simplifying t = -3.2 this is the right answer

OpenStudy (anonymous):

thaats what i said

OpenStudy (anonymous):

-3.2

OpenStudy (jordanloveangel):

oh my bad i read it wrong

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

its okay

OpenStudy (anonymous):

thank you xD

OpenStudy (jordanloveangel):

np

OpenStudy (anonymous):

100% mastered!

OpenStudy (jordanloveangel):

cool

OpenStudy (anonymous):

ty ty ty

OpenStudy (jordanloveangel):

np np np